# Evolution with Opponent-Learning Awareness

### v3

@@ Abstract.
The universe involves many independent co-learning agents as an ever-evolving part of our observed environment. Yet, in practice, Multi-Agent Reinforcement Learning (MARL) applications are usually constrained to small, homogeneous populations and remain computationally intensive. In this paper, we study how large heterogeneous populations of learning agents evolve in normal-form games. We show how, under assumptions commonly made in the multi-armed bandit literature, Multi-Agent Policy Gradient closely resembles the Replicator Dynamic, and we further derive a fast, parallelizable implementation of Opponent-Learning Awareness tailored for evolutionary simulations. This enables us to simulate the evolution of very large populations made of heterogeneous co-learning agents, under both naive and advanced learning strategies. We demonstrate our approach in simulations of 200,000 agents, evolving in the classic games of Hawk-Dove, Stag-Hunt, and Rock-Paper-Scissors. Each game highlights distinct ways in which Opponent-Learning Awareness affects evolution.
Key words and phrases:
Multi-agent reinforcement learning, Evolutionary game theory, Simulation, Opponent-learning awareness
1. Introduction
While most Reinforcement Learning (RL) theory relies on the stationarity assumption Sutton (2018), real-world situations are hardly ever truly stationary. Our universe is a place of perpetual change, where countless agents co-exist, co-learn and co-evolve. For any individual agent, other learning agents represent a fundamentally non-stationary part of the environment, and even more so when incentives are in conflict. In the realm of Reinforcement Learning, the study of this type of possibly adversarial non-stationarity Papoudakis et al. (2019) is a field known as Multi-Agent Reinforcement Learning (MARL).
While some notable achievements do tackle MARL scenarios, these are largely constrained by the high inherent complexity of multi-agent training. Successful MARL applications rarely feature more than a few agents, and in fact often resort to training a single neural network against one or few copies of itself. This powerful learning trick, named “self-play”, was famously used to learn Go at a superhuman level Silver et al. (2016), and competitive human-level team play in a video game Berner et al. (2019). On both occasions, the authors had to spend extreme amounts of compute, even though the number of agents involved was small and despite the computational savings enabled by training a single replicated model (as opposed to independent agents). Furthermore, recent successful MARL achievements typically relied on said “Multi-Agent” versions of single-agent RL algorithms (MADDPG Lowe et al. (2020), MAPPO Yu et al. (2022)…) which fall under a category of MARL techniques broadly referred to as “centralized training” Wong et al. (123). Centralized training consists of making use of privileged global information during training to reduce the fundamental non-stationarity provoked by the evolution of others, for instance by feeding this information to a centralized critic.
In this paper, we are instead interested in studying naturally plausible evolution processes. While centralized approaches facilitate the training of deployable, high-performance policies, they of course do not qualify as such processes, nor does self-play. However, there exists a line of biologically-inspired work that is close in philosophy to self-play: “Population-Based Training” (PBT) Jaderberg et al. (2017). In PBT, a population of policies is trained and periodically evaluated in a series of tournaments. These tournaments designate the “fittest” few policies to “select” and “replicate” before training continues. PBT is loosely part of a class of genetically inspired algorithms known as Evolution Strategies (ES).
The selection and replication heuristics used in ES are inspired by “imitation of the fittest”, a gradient-free revision protocol through which populations of non-learning agents are thought to evolve in nature Apesteguia et al. (2007). When aggregated across many agents, imitation of the fittest yields the Replicator population dynamic Sandholm (2010), arguably the most famous equation in Evolutionary Game Theory (EGT) Weibull (1997). Whereas MARL is often practically constrained to few-agent scenarios, EGT has long studied the evolution of large populations made of independent non-learning agents. Because these agents are non-learning, the revision protocols considered in EGT are extrinsic, similar to the genetically-inspired replicating and pruning performed in ES. However, imitation of success can also be considered as a very simple intrinsic learning protocol, where the agent simply copies the policy of its opponent based on their payoff. In this work, we bring advanced intrinsic MARL protocols to the scale of evolutionary simulations. In particular, we are interested in the evolutionary effects of Opponent-Learning Awareness, an advanced MARL protocol able to take advantage of non-stationary dynamics Foerster et al. (2017). Our contribution is threefold:
  • We show how, under common assumptions from the RL bandit literature, naive multi-agent Policy Gradient mathematically aligns with the Replicator Dynamic,
  • From this result, we derive fast, parallelizable implementations of both multi-agent Policy Gradient and Opponent-Learning Awareness, targeting evolutionary simulations in pairwise matrix games,
  • We perform population simulations of 200,000 learning agents and demonstrate different effects of both naive learning and Opponent-Learning Awareness on evolutionary dynamics.
2. Preliminaries
2.1. Evolutionary Game Theory
@@
EGT models evolution as a series of random pairwise interactions, where interactions are typically simple bi-matrix games (i.e., 2-player multi-armed bandits). Agents are sampled from a large population to be randomly paired and evaluated against their drawn opponent. The outcome of this interaction is a payoff for each opponent, whose expectation is called the agent’s fitness against the current population. An agent’s fitness depends on both its policy and the current configuration of the population. The agent’s policy is called its type, which is one of $n$ possible types. Agents with a greater fitness replicate and thus tend to “invade” the population, whereas agents with a lower fitness tend to go “extinct”. In particular, EGT is interested in evolutionarily stable equilibria, which are configurations of the population where the different types are present in stable proportions under replication dynamics. In evolutionarily stable equilibria, the population configuration is robust to rare mutations, where few agents randomly switch from one type to another. In this paper, we will be studying more complex, learning-based types of evolutionary equilibria in three classic symmetric games: Stag Hunt, Hawk-Dove, and Rock-Paper-Scissors.
2.1.1. Stag hunt
Stag hunt (SH) is a 2-action game illustrating the evolution of cooperation. In this game, agents need to hunt for food and choose to either go for a Stag, or go for a Hare. Hunting a Hare is easy: any agent choosing this option successfully receives a payoff of 1111111We choose all payoff scales arbitrarily.. Hunting a Stag is harder: both agents need to cooperate, otherwise the agent choosing to go for a Stag fails to catch anything and receives a payoff of 00. However, if both agents cooperate, they succeed and each receives a payoff of $1.8$, which is better than going for Hares. The bi-matrix representing this game is:
$\begin{array}[]{r|cc}&\text{Stag}&\text{Hare}\\ \hline\cr\textbf{Stag}&1.8,1.8&0,1\\ \textbf{Hare}&1,0&1,1\\ \end{array}$
where rows represent the action chosen by the ego agent (bold) with corresponding payoffs in first position, and columns represent the action chosen by the other agent with corresponding payoffs in second position. The game of Stag hunt has two distinct pure strategy Nash equilibria2222-player equilibria where each agent’s best response to the other player’s strategy is to always deterministically select the same action. Nash et al. (1950): (1) both agents always playing Stag, and (2) both agents always playing Hare.
2.1.2. Hawk-Dove
Hawk-Dove (HD) is a 2-action game illustrating the evolution of conflict. Agents either choose to act as a “Hawk” or as a “Dove”. When a Dove encounters another Dove, they share the available food (each receives a payoff of $1$). When a Dove encounters a Hawk, it yields and gets no food (payoff of 00) while the Hawk gets all of it (payoff of $2$). But when a Hawk encounters another Hawk, they fight and both get injured (payoff of $-1$ each). The bi-matrix representing this game is:
$\begin{array}[]{r|cc}&\text{Hawk}&\text{Dove}\\ \hline\cr\textbf{Hawk}&-1,-1&2,0\\ \textbf{Dove}&0,2&1,1\\ \end{array}$
The game of Hawk-Dove has two pure strategy Nash equilibria: (1) Hawk-Dove and (2) Dove-Hawk. But note that in these equilibria, Doves have a smaller payoff than Hawks. In the context of EGT, this means that when Doves encounter almost only Hawks, they go extinct and Hawks invade. However, when Hawks encounter almost always Hawks, their payoff is even less than Doves encountering Hawks, and thus Hawks move toward extinction as Doves invade. In other words, there are population configurations in which it is not more relevant to be a Hawk than a Dove in terms of fitness, and evolutionary dynamics naturally drive the population there. Examples of such configurations are (1) a population where all agents pick Hawk or Dove uniformly at random (which is the mixed-strategy Nash equilibrium3332-player equilibrium where both agents assign probabilities to each action. of this game), (2) a population formed of half deterministic Hawks and half deterministic Doves, or (3) any arbitrary population whose average behavior is to choose Hawk and Dove uniformly at random (for instance, half agents with policy $(\frac{2}{3},\frac{1}{3})$ and half agents with policy $(\frac{1}{3},\frac{2}{3})$…).
2.1.3. Rock-Paper-Scissors
Rock-Paper-Scissors (RPS) is a 3-action zero-sum [^4] The sum of the two agents’ payoffs is always 00. game. It illustrates more complex situations where there are cycles in the preferences over actions. Scissors beats Paper, Paper beats Rock, and Rock beats Scissors. The bi-matrix representing this game is:
@@ 2.1.3. Rock-Paper-Scissors
RPS has a single mixed-Nash equilibrium, where both players choose their actions uniformly at random. Similarly to the HD game, all populations whose average behavior is this equilibrium are neutrally stable under replication dynamics and drifting due to random mutations.
2.1.4. Population vector
A population whose individuals are distributed amongst $n$ types can be represented as a population vector $P\in\mathbb{R}^{n}$:
$P=\left(\begin{array}[]{c}p_{1}\\ p_{2}\\ \vdots\\ p_{n}\end{array}\right)\begin{array}[]{l}\text{proportion of type }1\\ \text{}\\ \text{}\\ \text{proportion of type }n\\ \end{array}$ (1)
2.1.5. Replicator Dynamic
Under the “imitation of the fittest” revision protocol (as well as several other bio-inspired revision protocols), large populations are known to follow a famous population dynamic over time ($t$), called the Replicator Dynamic:
@@ 2.1.5. Replicator Dynamic
2.2. Multi-Agent Reinforcement Learning
Whereas EGT typically studies evolution in the context of non-learning agents driven by extrinsic replication protocols, our work introduces intrinsic Multi-Agent Reinforcement Learning to the discussion. Instead of studying the evolutionary effects of what would be the selection heuristic of PBT, we study the evolutionary effects of what would be its learning process. Similar to EGT, we model evolution as a series of random pairwise interactions. Therefore, we are principally interested in 2-agent learning rules. In this paper, we will be specifically looking at two important such learning rules: policy gradient (PG), also referred to as “naive learning” in the MARL literature, and learning with opponent-learning awareness (LOLA) Foerster et al. (2017) in its full form (that is, using all terms in the first-order Taylor expansion, as opposed to Foerster et al. (2017)).
2.2.1. Policy gradient
@@
2.2.2. Opponent-learning awareness
Learning with opponent learning awareness (LOLA) Foerster et al. (2017) is an improved version of PG that takes into account the learning process of the other agent. More precisely, LOLA models the learning process of agent $2$ as if agent $2$ were a naive learner and differentiates through its learning step:
@@ 2.2.2. Opponent-learning awareness
Differentiating through the learning step of the opponent has two important advantages in our discussion: (1) it is a naturally plausible way of predicting the non-stationarity of the other agent (assuming agents maintain an internal model of others) in order to adapt beforehand, and (2) it enables the ego agent to adapt its own updates in order to steer the learning updates of the other agent toward its own incentives.
2.3. Population-policy equivalence
The population-policy equivalence noted by Bloembergen et al. Bloembergen et al. (2015) provides an interesting perspective binding evolution and Reinforcement Learning together. Let the population vector $P$ from Equation 1 represent a population of agents playing only pure strategies, i.e., where $p_{i}$ is the proportion of agents always playing action $i$. Uniformly sampling agents from this population is equivalent to sampling actions from the abstract stochastic policy defined by the probability vector $P$. Thus, Equation 3 can be viewed as a learning process, albeit at the population level, where the evolving population of non-learning agents $P$ is itself a self-play learning agent. In fact, Bloembergen et al. Bloembergen et al. (2015) show that this learning process is Cross Learning Cross (1973). In Section 3, we show how it is possible to study large populations of learning agents with a similar analysis.
3. Methods
To model how learning affects evolution, we adopt a philosophy similar to EGT. Namely, we consider large populations of independent learning agents, which are paired randomly at each evolution iteration and interact in normal-form matrix games. Each agent has its own learning rule (i.e., either of the two presented in Section 2.2) that it applies to its own policy after each pairwise interaction.
Whereas these rules are usually thought of in the context of persistent interactions between fixed agent pairs, in the evolutionary setting, they instead get applied after single interactions between random pairs. In other words, learning agents consider that the random opponent they interact with at each evolution step is a representative sample of the population and, for LOLA, of the current direction of its non-stationary dynamics. As hinted in Section 2.1, this yields complex population dynamics that may or may not converge to homogeneous populations, depending on factors such as the game, the learning rule, and the policy architecture.
3.1. Policy architecture
From an RL perspective, the normal-form matrix games presented in Section 2.1 are 2-agent multi-armed bandits. As this is a common assumption in multi-armed bandits Sutton (2018), we consider the policy architecture parameterized by the preference vector $\theta\in\mathbb{R}^{n}$, where $n$ is the number of actions, projected to the probability simplex by a simple softmax function $\sigma$, which yields the probability vector $P\in\mathbb{R}^{n}$ of the policy selecting each of the $n$ available actions:
@@ 3.1. Policy architecture
3.2. Analytical Policy Gradient

### v4

# Evolution of Societies via Reinforcement Learning

@@ Abstract
The universe involves many independent co-learning agents as an ever-evolving part of our observed environment. Yet, in practice, Multi-Agent Reinforcement Learning (MARL) applications are typically constrained to small, homogeneous populations and remain computationally intensive. We propose a methodology that enables simulating populations of Reinforcement Learning agents at evolutionary scale. More specifically, we derive a fast, parallelizable implementation of Policy Gradient (PG) and Opponent-Learning Awareness (LOLA), tailored for evolutionary simulations where agents undergo random pairwise interactions in stateless normal-form games. We demonstrate our approach by simulating the evolution of very large populations made of heterogeneous co-learning agents, under both naive and advanced learning strategies. In our experiments, 200,000 PG or LOLA agents evolve in the classic games of Hawk-Dove, Stag-Hunt, and Rock-Paper-Scissors. Each game provides distinct insights into how populations evolve under both naive and advanced MARL rules, including compelling ways in which Opponent-Learning Awareness affects social evolution.
(a) PG in SH ($s=1.8$)
(b) PG in HD ($f=-2$)
(c) PG in RPS
(d) LOLA in SH ($s=1.8$)
(e) LOLA in HD ($f=-2$)
(f) LOLA in RPS
Figure 1: Populations of 200,000 RL agents evolving in the classic games of Stag Hunt, Hawk-Dove and Rock-Paper-Scissors (columns), via Policy Gradient and LOLA (rows). Each agent is a stochastic policy, represented as linear coordinates between pure strategies. Dark shades of blue indicate high concentrations of agents, and evolution steps correspond to one learning step performed per agent. In Hawk-Dove, black dots indicate the average policy over the entire population.
(a) Stag Hunt
(b) Hawk Dove
Figure 2: Final average policy over the population, depending on cost values.
1 Introduction
Our universe is one of perpetual change, where countless agents co-exist, co-learn and co-evolve. From an individual agent’s perspective, other learning agents are a fundamentally non-stationary part of the environment, especially when incentives are in conflict (Papoudakis et al., 2019). In the realm of Reinforcement Learning (RL), the study of this type of possibly adversarial non-stationarity is a field known as Multi-Agent Reinforcement Learning (MARL). Some notable achievements do tackle MARL scenarios, but these are largely constrained by the high complexity of multi-agent training. In fact, these often resort to “self-play” (Silver et al., 2016; Berner et al., 2019), i.e., training a single neural network against one or few copies of itself, or to “centralized training” (Lowe et al., 2020; Yu et al., 2022), i.e., circumventing non-stationarity by using privileged global information.
We are interested in modeling plausible real-world social evolution processes that stem from RL revision protocols. We make the assumption that individuals continuously adapt their own strategies by following local RL rules, and we simulate how large societies evolve as a result. This setting is relevant to the study of social dynamics where individuals actively optimize their own fitness, such as economy and finance. More precisely, the scope of this paper is to simulate large populations of persistent RL agents, following an interaction model close to that of Evolutionary Game Theory (EGT). At each evolution step, agents are paired according to some assortment process (in our case, uniformly at random), and interact according to their respective policies. After each such interaction, agents adapt their policies according to their respective payoff and learning rule [^1] Throughout this paper, we use vocabulary from RL and EGT interchangeably. In particular, “payoff” = “return”, “fitness” = “value”, “strategy” = “policy”, “pure strategy” = “action”, “revision protocol” = “learning rule”.. Note that this type of social evolution is intrinsically driven by learning and individual preferences, in stark contrast to genetic evolution, which the literature usually considers as an extrinsic process happening among non-learning agents through their replication and death (Weibull, 1997). In our work, the concept of fitness does not necessarily refer to a measure of reproductive ability. Instead, it refers to a measure of expected completion of an arbitrary individual objective, i.e., an RL value.
We bring advanced intrinsic MARL protocols to the scale of evolutionary simulations, studying in particular the evolutionary effects of Opponent-Learning Awareness, an advanced MARL protocol able to take advantage of non-stationary dynamics (Foerster et al., 2017).
2 Preliminaries
This section reviews the concepts from Evolutionary Game Theory (EGT) and Multi-Agent Reinforcement Learning (MARL) used in this paper. A high-level overview of the literature contextualizing our work is provided in the Supplementary Material (Appendix A).
2.1 Evolutionary Game Theory (EGT)
@@
EGT models evolution as a series of random pairwise interactions, where interactions are typically simple bi-matrix games (i.e., 2-player multi-armed bandits). Agents are sampled from a large population to be randomly paired and evaluated against their drawn opponent. The outcome of this interaction is a payoff for each opponent, whose expectation is called the agent’s fitness against the current population. An agent’s fitness depends on both its policy and the current configuration of the population. The agent’s policy is called its type, which is one of $n$ possible types. In gene-inspired revision protocols, agents with a greater fitness replicate and thus tend to “invade” the population, whereas agents with a lower fitness tend to go “extinct”. In particular, EGT is interested in evolutionarily stable equilibria, which are configurations of the population where the different types are present in stable proportions under replication dynamics. In evolutionarily stable equilibria, the population configuration is robust to rare mutations, where few agents randomly switch from one type to another. In this paper, we will be studying more complex, learning-based evolutionary equilibria in three classic symmetric games: Stag Hunt, Hawk-Dove, and Rock-Paper-Scissors. These games can be described by bi-matrices:
$\begin{array}[]{r|cc}&\text{Stag}&\text{Hare}\\ \hline\cr\textbf{Stag}&s,s&0,1\\ \textbf{Hare}&1,0&1,1\\ \end{array}$
(a) Stag Hunt
$\begin{array}[]{r|cc}&\text{Hawk}&\text{Dove}\\ \hline\cr\textbf{Hawk}&f,f&2,0\\ \textbf{Dove}&0,2&1,1\\ \end{array}$
(b) Hawk-Dove
@@
(c) Rock-Paper-Scissors
where rows represent the action chosen by the ego agent (bold) with corresponding payoffs in first position, and columns represent the action chosen by the other agent with corresponding payoffs in second position.
Stag hunt (SH) is a 2-action game illustrating the evolution of cooperation. In this game, agents need to hunt for food and choose to either go for a Stag, or go for a Hare. Hunting a Hare is easy: any agent choosing this option successfully receives a payoff of $1$. Hunting a Stag is harder: both agents need to cooperate, otherwise the agent choosing to go for a Stag fails to catch anything and receives a payoff of 00. However, if both agents cooperate, they succeed and each receives a payoff of $s>1$, which is better than going for Hares. The game of Stag hunt has two distinct pure strategy Nash equilibria2222-player equilibria where each agent always selects the same action.: (1) both agents always playing Stag, and (2) both agents always playing Hare.
Hawk-Dove (HD) is a 2-action game illustrating the evolution of conflict over shareable resources. Agents either choose to act as a “Hawk” or as a “Dove”. When a Dove encounters another Dove, they share the available food (each receives a payoff of $1$). When a Dove encounters a Hawk, it yields and gets no food (payoff of 00) while the Hawk gets all of it (payoff of $2$). But when a Hawk encounters another Hawk, they fight and both get injured (payoff of $f<0$). The game of Hawk-Dove has two pure strategy Nash equilibria: (1) Hawk-Dove and (2) Dove-Hawk. But note that in these equilibria, Doves have a smaller payoff than Hawks. In the context of evolutionary genetics, this means that when Doves encounter almost only Hawks, they move toward extinction as Hawks invade. However, when Hawks encounter almost always Hawks, their expected payoff is even less than Doves encountering Hawks, and thus Hawks move toward extinction as Doves invade. In other words, there are population configurations in which it is not more relevant to be a Hawk than a Dove in terms of fitness, and replication dynamics naturally drive the population there.
Rock-Paper-Scissors (RPS) is a 3-action zero-sum [^3] The sum of the two agents’ payoffs is always 00. game. It illustrates more complex situations where there are cycles in the preferences over actions. Scissors beats Paper, Paper beats Rock, and Rock beats Scissors. RPS has a single mixed-Nash equilibrium, where both players choose their actions uniformly at random. Similarly to the HD game, all populations whose average behavior is this equilibrium are neutrally stable under replication dynamics and drifting due to random mutations.
A population whose individuals are distributed amongst $n$ types can be represented as a population vector $P\in\mathbb{R}^{n}$ whose components $0\leq p_{i}\leq 1$ sum to $1$ and represent the proportion of type $i$. Under the “imitation of the fittest” revision protocol (as well as several other bio-inspired revision protocols), large populations are known to follow a famous population dynamic over time ($t$), called the Replicator Dynamic:
@@
2.2 Multi-Agent Reinforcement Learning
While EGT traditionally focuses on populations where non-learning agents reproduce based on their fitness, our work examines how populations evolve when agents actively learn and adapt their strategies. Similar to EGT, we model social dynamics as a series of random pairwise interactions. Therefore, we are principally interested in 2-agent learning rules. In this paper, we will be specifically looking at two important such learning rules: policy gradient (PG), also referred to as “naive learning” in the MARL literature, and learning with opponent-learning awareness (LOLA) in its full form (i.e., using all terms from the first-order Taylor expansion).
@@ 2.2 Multi-Agent Reinforcement Learning
Learning with opponent learning awareness (LOLA) is an improved version of PG that takes into account the learning process of the other agent. More precisely, LOLA models the learning process of agent $2$ as if agent $2$ were a naive learner and differentiates through its learning step:
@@
Differentiating through the learning step of the opponent has an important advantage in our discussion: it is a naturally plausible way of predicting non-stationarity (assuming we maintain an internal model of others) in order to adapt beforehand and actively steer it toward our own incentives.
2.3 Population-policy equivalence
In a population of agents playing only pure strategies, uniformly sampling agents is equivalent to sampling actions from the abstract stochastic policy defined by the probability vector $P$. Thus, Equation 2 can be viewed as a learning process, albeit at the population level, where the evolving population of non-learning agents $P$ is itself a self-play learning agent (Bloembergen et al., 2015).
3 Methods
To model how learning affects societies, we adopt a philosophy similar to EGT. Namely, we consider large populations of independent learning agents, which are paired randomly at each evolution iteration and interact in normal-form matrix games. Each agent has its own learning rule (i.e., either of the two presented in Section 2.2) that it applies to its own policy after each pairwise interaction. Whereas MARL usually thinks about these rules in the context of persistent interactions between fixed pairs of agents, in the evolutionary setting, they instead get applied after single interactions between random pairs. In other words, learning agents consider that the random opponent they interact with at each evolution step is a representative sample of the population and, for LOLA, of the current direction of its non-stationary dynamics.
3.1 Policy architecture
From an RL perspective, the normal-form matrix games presented in Section 2.1 are 2-agent multi-armed bandits. As this is a common assumption in multi-armed bandits (Sutton, 2018), we consider the policy architecture parameterized by the preference vector $\theta\in\mathbb{R}^{n}$, where $n$ is the number of actions, projected to the probability simplex by a simple softmax function $\sigma$, which yields the probability vector $P\in\mathbb{R}^{n}$ of the policy selecting each of the $n$ available actions:
@@ 3.1 Policy architecture
3.2 Analytical Policy Gradient